# Barbie: Text to Barbie-Style 3D Avatars

### v6

@@ III-F Implementation Details
Our algorithm is implemented using PyTorch [56] and ThreeStudio [57]. We use MLP with multi-resolution hash encoding [49] to efficiently represent SDF, mSDF, and color fields. Given a text describing a dressed human, e.g., “A man wearing X1 and X2”, we first generate a base human body based on “A man in his underwear”. We then generate the apparel X1 based on “A piece of X1” and “A man wearing X1”, and then generate the apparel X2 based on “A piece of X2” and “A man wearing X1 and X2”. Finally, fine-tune the overall texture based on the full input text “A man wearing X1 and X2”. Additional details can be found in Supp. Mat.
@@ IV-A Experimental Settings
Baselines. We perform both quantitative and qualitative comparisons of Barbie with state-of-the-art (SOTA) SDS-based methods for text-driven avatar generation and apparel generation. For avatar generation, we compare with text-to-holistic-avatar methods (DreamWaltz [31], TADA [23], X-Oscar [36], HumanGaussian [32], HumanNorm [34], and DreamWaltz-G [64]) and text-to-decoupled-avatar method (SO-SMPL [35]). For text-to-apparel generation, we compare with text-to-3D methods (MVDream [16], GaussianDreamer [19], RichDreamer [18], LucidDreamer [21]), text-to-garment method (GarmentDreamer [39]), and SO-SMPL [35].
@@
DreamWaltz 0.5819 0.5333 3.67 3.67
TADA 0.5306 0.5333 4.67 5.00
X-Oscar 0.5069 0.5292 4.00 4.00
HumanGaussian 0.6222 0.5069 3.67 3.00
HumanNorm 0.5611 0.5292 5.00 3.33
DreamWaltz-G 0.6444 0.5722 2.33 3.33
SO-SMPL 0.6500 0.6000 0.00 0.67
Barbie (Ours) 0.7167/0.7333 0.6333/0.7000 76.67 77.00
MVDream 0.7000 0.5933 12.33 12.00
GaussianDreamer 0.7017 0.5283 5.00 5.33
RichDreamer 0.8100 0.6667 2.67 5.00
LucidDreamer 0.7533 0.6400 6.67 5.33
GarmentDreamer 0.6500 0.4500 1.00 1.00
SO-SMPL 0.8500 0.7667 0.00 0.00
Barbie (Ours) 0.8667/0.9667 0.7667/0.9167 72.33 71.33
  • The best result is highlighted in bolded. GQP: generation quality preference (%), TAP: text-image alignment preference (%). $\dagger$: When calculating VQA metrics, we ignore the problem of retrieving shoes and accessories, as GarmentDreamer [39] and SO-SMPL [35] not only support clothing generation.
Dataset Construction. We utilized ChatGPT [1] to randomly generate 30 text descriptions of dressed avatars, with each example wearing a top, a bottom, a pair of shoes, and two random accessories. 30 descriptions for dressed humans are used to evaluate avatar generation, and 30$\times$5 apparel descriptions are used to evaluate apparel generation (see the Supp. Mat. for the complete descriptions).
@@
Consequently, inspired by Progressive3D [55], we adopt fine-grained text-to-image evaluation metrics, including BLIP-VQA [59, 58] and BLIP2-VQA [60, 58], to evaluate the generation capacity of current methods and Barbie. Specifically, we first convert the prompt into multiple separate questions to retrieve corresponding content, then feed the rendered image of the generated content into the VQA model and ask questions one by one, and finally use the probability of answering “yes” as the evaluation metric. For instance, the input avatar prompt “A man wearing X1.” is converted into “Is the person in the picture a man?” and “Is the person in the picture wearing X1?”. The input apparel prompt “A pair of X1.” is converted into “Is the object in the picture a pair of X1?”. Besides, we randomly select 10 examples from the generated results to conduct a user study and ask 30 volunteers to assess (1) generation quality and (2) text-image alignment, and select the preferred methods.
@@ IV-B Comparisons of Avatar Generation
Comparison with Text-to-Avatar Methods. As shown in the upper part of Table II, our method significantly outperforms all baseline methods across all evaluation metrics. The qualitative results presented in the left part of Fig. 6 further validate its superiority. Compared to existing methods, Barbie generates avatars with more detailed and realistic geometry, while achieving better alignment with the input text, without omitting any specified apparel items. Moreover, Barbie achieves a high degree of disentanglement among body, garments, shoes, and accessories, enabling flexible outfit combinations and edits, much like dressing physical Barbie dolls (Fig. 9-(a) and -(b)). Furthermore, thanks to the use of G-Shell and the proposed SMPLX-evolving prior loss, Barbie supports expressive animation and compatibility with physical simulation (Fig. 9-(c) and -(d)).
Comparison with Text-to-Decoupled-Avatar Method. As illustrated in Fig. 7, our method surpasses SO-SMPL [35] in generating finer geometric details and more lifelike textures. Additionally, Barbie enables the generation of various accessories (e.g., necklaces, glasses, and watches), which are not supported by SO-SMPL [35].
Figure 8: Ablation Study of Barbie.
@@ IV-C Comparisons of Apparel Generation
As shown in the lower part of Table II, our approach consistently outperforms all competing methods on all evaluation metrics. Qualitative comparisons are shown in the right part of Fig. 6, displaying representative generated outfits. Compared to other methods, Barbie produces apparel with higher-fidelity geometry and textures, while maintaining accurate alignment with the input text, without introducing irrelevant elements such as body parts. Thanks to the expressive G-Shell [63] representation, we are able to accurately model the non-watertight topology of garments, enabling compatibility with physical simulations (Fig. 9-(d)).
@@ IV-D Ablation Study
We conduct detailed ablation studies in Fig. 8 to validate the effectiveness of each component in Barbie.
@@
Effect of Object-Specific Diffusion Models. To enhance in-domain outfit fidelity, we employ object-specific diffusion models for generating apparel geometry and texture. As illustrated in Fig. 8-(b), object-specific generative priors significantly enhance the visual quality of generated outfits, particularly in geometric detail.
@@
Figure 9: Applications of Barbie.
@@ IV-E Applications
As shown in Fig. 9-(a) and -(b), Barbie’s fine-grained decoupling enables the free combination and editing of any human body and outfit, similar to dressing physical Barbie dolls. This significantly enhances the playability and reusability of text-to-avatar generation. Thanks to the SMPLX-evolving prior loss $\mathcal{L}_{prior}$, we achieve accurate alignment between the generated assets and the SMPL-X model, enabling natural full-body animation involving facial expressions, body poses, and hand gestures, as demonstrated in Fig. 9-(c). The expression sequences are taken from TalkShow [92], and the body motions from AIST++ [93]. Furthermore, leveraging the expressive G-Shell [63] representation, our method accurately captures non-watertight garment topology, supporting physical simulation [91] as shown in Fig. 9-(d).
@@
Limited by the resolution of G-Shell, our method struggles to accurately model highly complex geometries such as hair, ears, and other facial features, as shown in Fig. 10. Leveraging recent advances in 3D representations [97] provides a promising path toward improving the quality of generated digital humans. Furthermore, our reliance on SDS loss introduces significant computational overhead, restricting practical application. Especially for decoupled avatar generation, the training time increases linearly according to the number of components. Future work will focus on integrating diverse human datasets and advanced priors to enable efficient, high-quality 3D Barbie-style avatar generation, comparable to feed-forward methods [94].
@@ VI Conclusion
We propose Barbie, a novel text-guided framework for creating an animatable 3D avatar dressed in decoupled shoes and accessories, along with simulation-ready garments, resembling iconic Barbie dolls. Specifically, we employ an expressive G-Shell to uniformly represent both watertight and non-watertight meshes. Additionally, we introduce an efficient mSDF initialization and a hole-preserving loss to ensure well-defined open surface modeling. To guarantee the domain-specific fidelity of the generated human body and outfits, we suitably combine different specific expert T2I models for domain-specific knowledge. To mitigate the negative impacts caused by the over-strong generative priors of the expert models, we propose a series of solid regularization losses and optimization strategies to ensure geometric rationality and texture harmony of the generated results. Extensive experiments demonstrate that our approach not only outperforms compared methods in both dressed avatar and apparel generation tasks, but also facilitates seamless composition and editing of outfits, as well as expressive animation and physical simulation.

### v7

@@ III-F Implementation Details
Our algorithm is implemented using PyTorch [50] and ThreeStudio [17]. Generating a human body takes $\sim$3 hours (24GB VRAM); generating an outfit takes $\sim$4 hours (40GB VRAM). We use MLP with multi-resolution hash encoding [47] to efficiently represent SDF, mSDF, and color fields. Given a text describing a dressed human, e.g., “A man wearing jeans and a T-shirt”, we first use a rule-based script to decompose the full input text into multiple sub-prompts: “A man in his underwear”, “A pair of jeans”, and “A T-shirt”. We then generate a base human body based on “A man in his underwear”. Next, we generate jeans based on “A pair of jeans” and “A man wearing jeans”, followed by generating the T-shirt based on “A T-shirt” and “A man wearing jeans and a T-shirt”. Finally, fine-tune the texture of the assembled avatar based on the full input text “A man wearing jeans and a T-shirt”. Additional details can be found in the Supp. Mat.
@@ IV-A Experimental Settings
Baselines. We perform both quantitative and qualitative comparisons of DreamBarbie with state-of-the-art (SOTA) SDS-based methods for text-driven avatar generation and apparel generation. For avatar generation, we compare with text-to-holistic-avatar methods (DreamWaltz [23], TADA [39], X-Oscar [45], HumanGaussian [41], HumanNorm [22], and DreamWaltz-G [24]) and text-to-decoupled-avatar method (SO-SMPL [66]). For text-to-apparel generation, we compare with text-to-3D methods (MVDream [62], GaussianDreamer [75], RichDreamer [54], LucidDreamer [38]), text-to-garment method (GarmentDreamer [31]), and SO-SMPL [66].
@@
DreamWaltz 0.5819 0.5333 2.80 2.31
TADA 0.5306 0.5333 3.25 2.91
X-Oscar 0.5069 0.5292 3.43 3.62
HumanGaussian 0.6222 0.5069 2.35 2.05
HumanNorm 0.5611 0.5292 3.81 3.51
DreamWaltz-G 0.6444 0.5722 2.87 2.54
SO-SMPL 0.6500 0.6000 0.45 0.71
DreamBarbie (Ours) 0.7167/0.7333 0.6333/0.7000 81.04 82.35
MVDream 0.7000 0.5933 9.03 8.92
GaussianDreamer 0.7017 0.5283 3.84 4.14
RichDreamer 0.8100 0.6667 3.25 2.95
LucidDreamer 0.7533 0.6400 4.59 3.77
GarmentDreamer 0.6500 0.4500 0.78 0.63
SO-SMPL 0.8500 0.7667 0.26 0.37
DreamBarbie (Ours) 0.8667/0.9667 0.7667/0.9167 78.25 79.22
  • The best result is highlighted in bolded. GQP: generation quality preference (%), TAP: text-image alignment preference (%). $\dagger$: When calculating VQA metrics, we ignore the problem of retrieving shoes and accessories, as GarmentDreamer [31] and SO-SMPL [66] only support clothing generation.
[…]
Figure 8: Ablation Study of DreamBarbie.
@@
Consequently, inspired by Progressive3D [9], we adopt fine-grained text-to-image evaluation metrics, including BLIP-VQA [34, 32] and BLIP2-VQA [33, 32], to evaluate the generation capacity of current methods and DreamBarbie. Specifically, we first convert the prompt into multiple separate questions to retrieve corresponding content, then feed the rendered image of the generated content into the VQA model and ask questions one by one, and finally use the probability of answering “yes” as the evaluation metric. For instance, the input avatar prompt “A man wearing X1.” is converted into “Is the person in the picture a man?” and “Is the person in the picture wearing X1?”. The input apparel prompt “A pair of X1.” is converted into “Is the object in the picture a pair of X1?”. Besides, we randomly select 40 cases (20 for avatar and 20 for outfit) to conduct a user study and ask 67 volunteers to assess (1) generation quality and (2) text-image alignment, and select the preferred methods.
@@ IV-B Comparisons of Avatar Generation
Comparison with Text-to-Avatar Methods. As shown in the upper part of Table II, our method significantly outperforms all baseline methods across all evaluation metrics. The qualitative results presented in the left part of Fig. 6 further validate its superiority. Compared to existing methods, DreamBarbie generates avatars with more detailed and realistic geometry, while achieving better alignment with the input text, without omitting any specified apparel items. Moreover, DreamBarbie achieves a high degree of disentanglement among body, garments, shoes, and accessories, enabling flexible outfit combinations and edits, much like dressing physical Barbie dolls (Fig. 9-(a) and -(b)). Furthermore, thanks to the use of G-Shell and the proposed SMPLX-evolving prior loss, DreamBarbie supports expressive animation and compatibility with physical simulation (Fig. 9-(c) and -(d)).
Comparison with Text-to-Decoupled-Avatar Method. As illustrated in Fig. 7, our method surpasses SO-SMPL [66] in generating finer geometric details and more lifelike textures. Additionally, DreamBarbie enables the generation of various accessories (e.g., necklaces, glasses, and watches), which are not supported by SO-SMPL [66].
@@ IV-C Comparisons of Apparel Generation
As shown in the lower part of Table II, our approach consistently outperforms all competing methods on all evaluation metrics. Qualitative comparisons are shown in the right part of Fig. 6, displaying representative generated outfits. Compared to other methods, DreamBarbie produces apparel with higher-fidelity geometry and textures, while maintaining accurate alignment with the input text, without introducing irrelevant elements such as body parts. Thanks to the expressive G-Shell [42] representation, we are able to accurately model the non-watertight topology of garments, enabling compatibility with physical simulations (Fig. 9-(d)).
@@ IV-D Ablation Study
We conduct detailed ablation studies in Fig. 8 to validate the effectiveness of each component in DreamBarbie.
@@
Effect of Dual-Space Optimization. To generate fine-detailed, well-fitted apparel, we employ dual-space optimization with human- and object-specific generative priors. As shown in Fig. 8-(b), human-space-only optimization yields assets with poor detail (orange box) and erroneously absorbs body parts into clothing (blue box), breaking semantic boundaries. In contrast, apparel-space-only optimization produces ill-fitting outfits—evident in oversized shoes (green box) and texture discontinuities near the waist (purple box)—undermining realism. Our dual-space approach jointly leverages both priors to achieve detailed, body-conforming garments and visually coherent avatars.
@@
Figure 9: Applications of DreamBarbie.
@@ IV-E Applications
As shown in Fig. 9-(a), DreamBarbie’s fine-grained decoupling enables the flexible combination of any human body with any outfit, similar to dressing physical Barbie dolls. Thanks to the SMPLX-evolving prior loss $\mathcal{L}_{prior}$, we achieve accurate alignment between the generated assets and the SMPL-X model. This alignment allows for intuitive apparel editing by adjusting shape parameters, and supports natural full-body animation through control of pose and expression parameters, as shown in Fig. 9-(b) and -(c). The expression sequences and body motions are taken from AIST++ [35] and TalkShow [74]. Furthermore, leveraging the expressive G-Shell [42] representation, our method accurately captures non-watertight garment topology, supporting physical simulation [63] as shown in Fig. 9-(d). Our method enables the creation of Barbie-style 3D avatar meshes. As a result, users can view and combine human bodies and outfits within professional 3D software [25], as illustrated in Fig. 9-(e).
@@
Limited by the resolution of G-Shell, our method struggles to accurately model highly complex geometries such as hair, ears, and other facial features, as shown in Fig. 10. In addition, while the use of predefined templates ensures reasonable initialization, it may slightly limit generation diversity. In future work, we aim to explore template-free, disentangled generation by leveraging 3D generative priors [72]. Finally, our reliance on the SDS loss introduces significant computational overhead, limiting practical applicability. As a direction for future work, we plan to adopt a feed-forward, disentangled generation framework such as StdGen [18], integrating diverse human datasets and advanced priors to enable efficient 3D Barbie-style avatar generation.
@@ VI Conclusion
We propose DreamBarbie, a novel text-guided framework for creating an animatable 3D avatar dressed in decoupled shoes and accessories, along with simulation-ready garments, resembling iconic Barbie dolls. Specifically, we employ an expressive G-Shell to uniformly represent both watertight and non-watertight meshes. We introduce an SDF-based initialization that reformulates boundary definition from manifold geodesics into 3D field intersections, significantly enhancing stability and efficiency of open-surface modeling. To guarantee the domain-specific fidelity of the generated human body and outfits, we suitably combine different specific expert T2I models for domain-specific knowledge. To mitigate the negative impacts caused by the over-strong generative priors of the expert models, we propose a series of solid regularization losses and optimization strategies to ensure geometric rationality and texture harmony of the generated results. Extensive experiments demonstrate that our approach not only outperforms compared methods in both dressed avatar and apparel generation tasks, but also facilitates seamless composition and editing of outfits, as well as expressive animation and physical simulation.